Wednesday, October 11, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #


  1. Commissioning plan updates from detector systems
  2. Status update from detector owners, calibrations, efficiencies, etc.
  3. Review systems owners and update information
  4. Address system specific points of interest for the Fall 2017 run



  • We briefly discussed the Fall 2017 Commissioning Plan
  • The summary table needs to be updated and tasks need to be assigned
  • We discussed the shift schedule
  • Shift crew will need to refer to a good wiki page
  • Dave G. Noted that ALL shift crew need to have the Hall walk through prior to locking up so that they can sit shift
  • We discussed the shift schedule
    • The official version and sign up period will go out in about a week or so
  • We need to ensure that we have the necessary RSAD and OSP documents readily available on the wiki
  • We need and OSP for the GEM and drift chambers


12 GeV Drift Chambers

  • Carlos showed a few log entries regarding the chamber work on the HMS platform
  • Chamber signals
    • Individual wire signals were investigated with a wire terminator and they look great
    • The logic pulse produced by the discriminator cards looks extremely stable
  • Wire maps were shown with cosmic data taken overnight
    • There are utilizing scintillator triggers with two on the top and two on the bottom
  • After fixing a connectivity issue the updated maps look good
  • Some efficiency plots were shown
    • They look rather good
    • 99% efficiency is expected in the range of 1900 - 2200 V which is currently being observed
    • The planes appear to plateau at 1940V at a threshold of 4.5 V


  • Simona will checkout the PMTs this afternoon
  • Only 3 planes will be checked out since 5 paddles from the first plane are being utilized for the 12 GeV chamber tests

Heavy Gas Cherenkov

  • C4F8O OR C4F10O will be in the tank
  • It is unclear at the moment which pressure will be utilized
    • This needs to be followed up on


  • Arshak reported that they plan to take cosmic data soon



  • Steve reported that the timing is now understood
  • The TI latency was being set to be extremely large
    • This was reduced and behaved in an expected way
  • The dynamic range in the timing now looks to be okay
  • The TI delay was reduced overall by approximately 0.5 us
  • The new optical fiber in SHMS hut is now being utilized instead of the previous HMS patch
  • Latif reported that overnight cosmic data has been taken
  • Pedestal correction and map work is ongoing
  • Work is ongoing to investigate the possibility of noise and zero suppression, Steve is looking into this
  • Latif is looking into how exactly to set the thresholds relative to the baseline

Noble Gas Cherenkov

  • Out for commissioning
  • Will be there for F2


  • Simona reported on the quartz plane
    • She can make 4 per day
    • 11 are curing currently and should be done by Saturday and then flipped and should be done by the following Saturday
  • All bars should be done around Nov 5
  • When the frame comes down S2X will come with it

Aerogel Cerenkov

  • Vlad reported they have a new HV module and it will be installed later this afternoon
    • The previous HV module has been causing problems
  • One PMT is questionable, but a spare is on the way
  • They are looking at data and it looks reasonable
  • This tube has been carefully investigated although they are almost certain it is the tube itself


  • They have been taking cosmic data and all channels are alive but will be looked at further